# Quantization-Free Autoregressive Action Transformer

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Generative modeling lies at the heart of modern machine learning, enabling the production of outputs that adhere to user-specified objectives and constraints by navigating high-dimensional spaces of potential outcomes. Fundamentally, this process can be understood as a form of search where naive or exhaustive strategies are computationally intractable due to the size of the space. We consider imitation learning for robotic systems as an extreme example, where the search space has uncountable cardinality and a complex structure arising from nonlinear dynamics, interactions with the environment and the fundamental multimodality in human or robot behavior.
In many real-world robotics applications, the control signals involved are continuous in nature (Muehlebach & D’Andrea, 2017; Zughaibi et al., 2024; Ma et al., 2023). Current state-of-the-art methods for modeling policies in continuous control tasks broadly fall into two categories: autoregressive transformer models (Lee et al., 2024; Shafiullah et al., 2022) and diffusion models (Chi et al., 2023; Wang et al., 2022; Psenka et al., 2023), which will be discussed in the next two paragraphs.
@@
Existing autoregressive policies, on the one hand, sidestep the challenge of learning in a continuous domain by discretizing the actions (Lee et al., 2024; Shafiullah et al., 2022). This discretization can introduce several drawbacks: It discards the inherent structure of the continuous space, increases complexity by adding a separate quantization step, and may limit expressiveness or accuracy when fine-grained control is required. Furthermore, the popular approach of learning a VQ-VAE (Van Den Oord et al., 2017) of the action space involves non-differentiable operations and hence requires advanced tricks during representation learning to side-step optimization difficulties and instabilities (Łańcucki et al., 2020; Kolesnikov et al., 2022; Huh et al., 2023; Mentzer et al., 2024). Hence, our goal is to design autoregressive transformers that preserve the continuous nature of the action space, leveraging the smooth and potentially multimodal behavior patterns that experts exhibit.
Figure 1: We present the diversity of behaviors generated by Q-FAT across three different environments. The first three images illustrate the initial task execution in the Franka Kitchen environment (Gupta et al., 2019), where the robot begins by either moving the kettle, opening the microwave, or interacting with the bottom burner. The third and fourth images showcase the UR3 BlockPush environment, demonstrating how the robotic manipulator initiates the task by pushing either the blue or red block. Finally, the last image depicts the PushT environment (Chi et al., 2023), where Q-FAT successfully captures the multimodal nature of the task, producing smooth and diverse motion trajectories.
@@
Diffusion policies, on the other hand, are adept at capturing continuous action distributions, but face two critical challenges in control tasks. First, real-time applications like robotics demand fast, efficient inference, yet diffusion models rely on iterative sampling, requiring dozens of forward passes to generate actions. This makes them orders-of-magnitude slower than autoregressive or feedforward policies (Lee et al., 2024). Second, the downstream tasks of imitation learning is often reinforcement learning (Schaal, 1996; Peters & Schaal, 2008; Zhu et al., 2018; Hu et al., 2024), which typically requires exact action likelihoods for tasks like exploration and off-policy corrections (Sutton et al., 1999; Abdolmaleki et al., 2018; Levine et al., 2020; Wu et al., 2019). Diffusion models, however, generate samples by integrating (stochastic) differential equations (Song et al., 2023, 2020), making likelihood estimation computationally expensive.
@@
To address these limitations, we propose the use of a Generative Infinite Vocabulary Transformer (GIVT) (Tschannen et al., 2024) that models the policy as a Gaussian Mixture Model (GMM) on top of a decoder-only transformer architecture (Vaswani, 2017; Radford et al., 2018). This approach enables modeling continuous actions as “tokens“ by directly parameterizing the components of the GMM in each time step, avoiding the need for discretization and thus preserving the inherent geometry of the action space. This choice allows us to exploit the benefits of large-scale sequence modeling, such as autoregressive policy generation and access to explicit likelihood estimates, while avoiding the pitfalls of artificially segmenting continuous control signals into discrete bins. In addition, we discuss two different sampling strategies that reduce variability in the generated trajectories.
@@
Figure 2: Overview of Q-FAT. A sequence of $h^{s}$ previous states and $h^{g}$ goal states are projected into the Transformer’s embedding dimension using a linear layer. The Transformer then predicts the action distribution by predicting the GMM means, variances and mixture probabilities.
@@ 2 Sequence to Action Prediction
We consider the following standard setting for imitation learning of simulated robotics tasks. An agent equipped with continuous controls $\mathcal{A}$ evolves in a continuous state space $\mathcal{S}$. In behavioral cloning, we assume that a so-called behavior policy $\pi_{b}$ collected a dataset $D=\{\tau_{i}\}_{i=0}^{N-1}$ of $N$ trajectories in the environment, where each trajectory $\tau_{i}=\{s^{i}_{0},a^{i}_{0},\dots,a^{i}_{n^{i}-1},s^{i}_{n^{i}}\}$ is a sequence of states $s^{i}_{t}\in\mathcal{S}$ and actions $a^{i}_{t}\in\mathcal{A}$, with $n^{i}$ denoting the length of trajectory $i$. Note that we do not impose a Markovian assumption on the behavior policy: in general $a^{i}_{t}$ might depend on all previous states $s^{i}_{<t}$.
@@
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=0}^{N-1}\sum_{t=0}^{n^{i}-1}\log\pi_{\theta}(a^{i}_{t}\mid s^{i}_{<t})\,.$ (1)
Note that a standard result in variational inference (Hoffman et al., 2013) shows that this objective is equivalent to minimizing the Kullback-Leibler (KL) divergence between the behavioral policy $\pi_{b}$ and the learned policy $\pi_{\theta}$, implying that the learnt policy $\pi_{\theta}$ should match, or clone, the behavior of $\pi_{b}$. However, the choice of parametrization is crucial to guarantee a good approximation of the behavior policy.
@@ 3 The Q-FAT Algorithm
Current transformer-based policy parametrization approaches (Lee et al., 2024; Shafiullah et al., 2022; Mete et al., 2024) consist of first discretizing the action space, e.g. using a VQ-VAE (Van Den Oord et al., 2017; Zeghidour et al., 2021), to comply with the standard transformer architecture requiring a discrete vocabulary. Inspired by the success of GIVT  (Tschannen et al., 2024) in the image generation domain, our model Quantization-Free Autoregressive Action Transformer (Q-FAT) adapts GIVT to the continuous action behavioral cloning setting, modeling the action distribution at each step by predicting the parameters of a multivariate Gaussian mixture model (GMM). Namely, Q-FAT models the probability of an action $a_{t}\in\mathbb{R}^{m}$ at step $t$ as
$\pi_{\theta}\bigl{(}a_{t}\,\big{|}\;s_{t-1},\dots,s_{t-h^{s}})\!=\!\sum_{i=1}^{k}\pi_{i}^{t}\,\mathcal{N}\!\Bigl{(}a_{t}\;\big{|}\;\boldsymbol{\mu}_{i}^{t},\boldsymbol{\sigma}_{i}^{t}\Bigr{)}\,,\quad$ (2)
where $\pi_{i}^{t}$, $\boldsymbol{\mu}_{i}^{t}$, and $\boldsymbol{\mathbf{\sigma}_{i}^{t}}$ are produced by the transformer decoder (Vaswani, 2017; Radford et al., 2018) at step $t$, conditioned on a history of previous states $\{s_{i}\}_{i=t-h^{s}}^{t-1}$ with $s_{i}\in\mathbb{R}^{d}$. Using a state history of length $h^{s}$ instead of only using the most recent state accounts for the potential non-Markovian structure in the demonstration dataset. Rather than predicting the full covariance $\mathbf{\Sigma}^{t}_{i}$ per mixture component, we assume a diagonal Gaussian distribution and only predict the diagonal entries of the covariance matrix $\boldsymbol{\sigma}^{t}_{i}$. While this simplifying assumption enables efficient likelihood computations, it does not impose independence across action dimensions in general.
@@ 3.1 Goal Conditioning
For some tasks such as tracking, a near-future goal can be appended to the input to condition the resulting policy. The conditioning is formalized by defining a set of goal states $\mathcal{G}\subseteq\mathcal{S}$. We then learn a goal-conditioned policy $\pi_{\theta}\bigl{(}a_{t}\,\big{|}\;s_{t-1},\dots,s_{t-h^{s}},g\bigr{)}$ where $g\subset G$. In practice, we define $g$ as a sequence of future states to be reached, i.e., $g=\{s^{g}_{i}\}_{i=t}^{t+h^{g}}$, where the goal states $s^{g}_{i}$ are typically generated from a high-level planner that plans instrumental goal states as done in Gupta et al. (2019). In our experiments, we simulate goal states from the demonstration dataset.
@@ 3.2 Action Sampling from a $k$-GMM
Q-FAT samples the next action from the conditional distribution, $a_{t}\sim\pi_{\theta}(\cdot|s_{t-1},\dots,s_{t-h^{s}},g)$, which is a GMM with $k$ components. In practice, the model can be trained to output a short sequence of actions that are then executed in an open-loop fashion. This is done to ensure action consistency and avoid over-fitting to idle actions (Chi et al., 2023). For simplicity, we describe only the next-action generation setting.
@@ 3.3 Stabilizing output trajectories
When controlling real-world robotics systems, it is important to generate action sequences without jitter, as high frequency inputs may damage actuators and excite undesirable structural modes (Astrom, 1995). Thanks to the richness of the GMM representation of the policy, we can directly design stabilization patches.
Figure 3: Potential detrimental effect of down-scaling the variance of the GMM mixture components when $k$ is misspecified. A uni-modal Gaussian distribution (gray) is approximated with two components (orange and green). The resulting density function (blue) fits the data distribution perfectly. Down-scaling the variance to $0.1$ (red) induces irreducible variance proportional to the distance between the modes.
@@
A common heuristic to mitigate noisy samples is to adjust the sampling temperature, which in our GMM parametrization is equivalent to down-scaling the variance in each Gaussian component (Mandlekar et al., 2021; Hoffman et al., 2020; Tschannen et al., 2024, 2025). Concretely, one replaces each covariance $\boldsymbol{\sigma}_{i}$ by $\alpha\,\boldsymbol{\sigma}_{i}$, for some small $\alpha\in(0,1)$—sometimes by several orders of magnitude ($\alpha\ll 1$). This compresses each component, forcing sampled actions to concentrate more tightly around each component’s mean. However, merely shrinking the variances does not address the large-scale spread when mixture components are far apart. The inter-component variance remains, and noisy samples may still arise from small but non-negligible transitions among disparate components. This phenomenon is exacerbated when the choice of the number of components $k$ is misspecified (see Figure 3 and further discussion in Appendix A). Furthermore, it has been shown by Carreira-Perpinan (2000) that GMMs in dimension $d\geq 2$ may have a mode that does not directly align with the mean of any component. In such a situation, the down-scaling technique is likely to result in an insidious loss of the extra mode (see Figure 4).
Mode-tracking via the Mean-Shift algorithm We explored a principled way to reduce the variance of sampling from multimodal distributions by explicitly identifying the major modes and then sampling from those modes with an appropriate probability while controlling variance. An effective way to find the modes in a GMM is via the mean-shift algorithm (Carreira-Perpinan, 2000), which is an efficient fixed-point iteration algorithm (see Appendix B for further details). We propose a multinomial distribution over the modes [^1] The number of modes $J$ is hard to control theoretically (Carreira-Perpinan, 2000). $\{m_{1},...m_{J}\}$ that respects the shape of the density:
@@
where $\text{where }w_{j}$ $=p_{\text{GMM}}\bigl{(}m_{j}\bigr{)}\times\bigl{|}-H_{j}\bigr{|}^{-\frac{1}{2}}$ (4)
and $\text{and }H_{j}$ $=\nabla^{2}\log p_{\text{GMM}}(m_{j})\,.$ (5)
@@
[…]
(a) Demonstration Dataset
(b) GMM Sampling
(c) $\sigma$-scaled Sampling
(d) Mode Sampling
Figure 5: We demonstrate the effect of multiple sampling techniques from a $8$-mixture Q-FAT policy trained on the toy multi-route environment (Shafiullah et al., 2022). (a) Visualization of the raw dataset, where the agent starts from the bottom-left corner and follows either the blue or the red path, with the goal of landing inside of the green target region on the top-right corner. Both paths are equally likely in the dataset. (b) Sampling directly from the GMM results in noisy samples because the model also captures and samples from the noise present in the dataset. (c) In $\sigma$-scaled sampling, the variances of the GMM components are scaled by a factor of $10^{-8}$. Despite reducing the variance per component, variance scaling does not decrease for the variance introduced due to the distance between the mixture components. (d) Mode sampling suppresses most of the noise in the demonstrations while preserving the multimodality of the original dataset.
@@ 4 Experiments
We evaluate the performance of Q-FAT through extensive experiments across four representative simulated robotics environments: PushT (Chi et al., 2023), Kitchen (Gupta et al., 2019), UR3 BlockPush (Florence et al., 2022) and Multimodal Ant (Lee et al., 2024). These environments cover the most common control signals in practice, namely position and velocity of robot’s end effector and robot joint angles. A detailed description of each environment is provided in Appendix C. Our experiments are designed to address the following key research questions:
@@
We find that Q-FAT achieves state-of-the-art performance across the majority of conditional and unconditional tasks, achieving state-of-the-art balance between the quality and diversity of the generated actions. We report the results of Q-FAT using the variance reduced sampling (Section 3.2) with a variance scaling of $10^{-6}$ and we discuss the implications of this choice in Section 4.3. A summary of our results is presented in Table 1.

### v3

@@ Abstract
footnotetext: Correspondence: Ziyad Sheebaelhamd <ziyad.sheebaelhamd@uni-tuebingen.de>
@@ 1 Introduction
Generative modeling lies at the heart of modern machine learning, enabling the production of outputs that adhere to user-specified objectives and constraints by navigating high-dimensional spaces of potential outcomes. Fundamentally, this process can be understood as a form of search where naive or exhaustive strategies are computationally intractable due to the size of the space. We consider imitation learning for robotic systems as an extreme example, where the search space has uncountable cardinality and a complex structure arising from nonlinear dynamics, interactions with the environment and the fundamental multimodality in human or robot behavior. In many real-world robotics applications, the control signals involved are continuous in nature [33, 59, 28]. Current state-of-the-art methods for modeling policies in continuous control tasks broadly fall into two categories: autoregressive transformer models [26, 45] and diffusion models [7, 54, 40], which will be discussed in the next two paragraphs.
@@
Existing autoregressive policies, on the one hand, ignore the challenge of learning in a continuous domain by discretizing actions [26, 45]. This discretization can introduce several drawbacks: It discards the inherent structure of the continuous space, increases complexity by adding a separate quantization step, and may limit expressiveness or accuracy when fine-grained control is required. Furthermore, the popular approach of learning a VQ-VAE [52] of the action space involves nondifferentiable operations and therefore requires advanced tricks during representation learning to sidestep optimization difficulties and instabilities [24, 20, 18, 31]. Hence, our goal is to design autoregressive transformers that preserve the continuous nature of the action space, leveraging the smooth and potentially multimodal behavior patterns that experts exhibit.
@@
Diffusion policies, on the other hand, are adept at capturing continuous action distributions, but face two critical challenges in control tasks. First, real-time applications like robotics demand fast, efficient inference, yet diffusion models rely on iterative sampling, requiring dozens of forward passes to generate actions. This makes them orders-of-magnitude slower than autoregressive or feedforward policies [26]. Second, the downstream task of imitation learning is often reinforcement learning [43, 39, 58, 17], which typically requires exact action likelihoods for tasks like exploration and off-policy corrections [49, 1, 27, 56]. Diffusion models, however, generate samples by integrating (stochastic) differential equations [48, 47], making likelihood estimation computationally expensive.
@@
To address these limitations, we propose the use of a Generative Infinite Vocabulary Transformer (GIVT) [50] that models the policy as a Gaussian Mixture Model (GMM) on top of a decoder-only transformer architecture [53, 41]. This approach enables modeling continuous actions as “tokens“ by directly parameterizing the components of the GMM in each time step, avoiding the need for discretization and thus preserving the inherent geometry of the action space. This choice allows us to exploit the benefits of large-scale sequence modeling, such as autoregressive policy generation and access to explicit likelihood estimates, while avoiding the pitfalls of artificially segmenting continuous control signals into discrete bins. In addition, we discuss two different sampling strategies that reduce variability in the generated trajectories.
@@
(a) An overview of the Q-FAT architecture.
(b) Sample trajectories generated by a Q-FAT policy.
Figure 1: Figure (a) shows the overview of QFAT. A sequence of $h^{s}$ previous states and $h^{g}$ goal states are projected into the Transformer’s embedding dimension using a linear layer. The Transformer then predicts the action distribution by predicting the GMM means, variances and mixture probabilities. Figure (b) shows a sample of generated trajectories from 3 representative environments, demonstrating the captured multi-modality in the sequence of solving tasks (Kitchen and UR3 Block Push) or the direction from which an object is approached (PushT).
@@ 2 Sequence to Action Prediction
We consider the following standard setting for imitation learning of simulated robotics tasks. An agent equipped with continuous controls $\mathcal{A}$ evolves in a continuous state space $\mathcal{S}$ in discrete time. In behavioral cloning, we assume that a so-called behavior policy $\pi_{b}$ collected a dataset $D=\{\tau_{i}\}_{i=0}^{N-1}$ of $N$ trajectories in the environment, where each trajectory $\tau_{i}=\{s^{i}_{0},a^{i}_{0},\dots,a^{i}_{n^{i}-1},s^{i}_{n^{i}}\}$ is a sequence of states $s^{i}_{t}\in\mathcal{S}$ and actions $a^{i}_{t}\in\mathcal{A}$, with $n^{i}$ denoting the length of trajectory $i$. Note that we do not impose a Markovian assumption on the behavior policy. In general $a^{i}_{t}$ might depend on a history of states $\{s^{i}_{j}\}_{j=t-h}^{t}$, where $h$ is a fixed window.
@@
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=0}^{N-1}\sum_{t=0}^{n^{i}-1}\log\pi_{\theta}(a^{i}_{t}\mid s^{i}_{t-h:t})\,.$ (1)
Note that a standard result in variational inference [15] shows that this objective is equivalent to minimizing the Kullback-Leibler (KL) divergence between the behavioral policy $\pi_{b}$ and the learned policy $\pi_{\theta}$, implying that the learned policy $\pi_{\theta}$ should match, or clone, the behavior of $\pi_{b}$. However, the choice of parametrization is crucial to guarantee a good approximation of the behavior policy.
@@ 3 The Q-FAT Algorithm
Current transformer-based policy parametrization approaches [26, 45, 32] consist of first discretizing the action space, e.g. using a VQ-VAE [52, 57], to comply with the standard transformer architecture requiring a discrete vocabulary. Inspired by the success of GIVT [50] in the image generation domain, our model Quantization-Free Autoregressive Action Transformer (Q-FAT) adapts GIVT to the continuous action behavioral cloning setting, modeling the action distribution at each step by predicting the parameters of a multivariate Gaussian mixture model (GMM). Namely, Q-FAT models the probability of an action $a_{t}\in\mathbb{R}^{m}$ at step $t$ as
$\pi_{\theta}\bigl(a_{t}\,\big|\;s_{t-1},\dots,s_{t-h^{s}})\!=\!\sum_{i=1}^{k}\pi_{i}^{t}\,\mathcal{N}\!\Bigl(a_{t}\;\big|\;\boldsymbol{\mu}_{i}^{t},\boldsymbol{\sigma}_{i}^{t}\Bigr)\,$$\quad$ (2)
where $\pi_{i}^{t}$, $\boldsymbol{\mu}_{i}^{t}$, and $\boldsymbol{\mathbf{\sigma}_{i}^{t}}$ are produced by the transformer decoder [53, 41] at step $t$, conditioned on a history of previous states $\{s_{i}\}_{i=t-h^{s}}^{t-1}$ with $s_{i}\in\mathbb{R}^{d}$. Using a state history of length $h^{s}$ instead of only using the most recent state accounts for the potential non-Markovian structure in the demonstration dataset. Rather than predicting the full covariance $\mathbf{\Sigma}^{t}_{i}$ per mixture component, we assume a diagonal Gaussian distribution and only predict the diagonal entries of the covariance matrix $\boldsymbol{\sigma}^{t}_{i}$. While this simplifying assumption enables efficient likelihood computations, it does not impose independence across action dimensions in general. Q-FAT is trained using standard teacher forcing with a causal attention mask. We highlight that we only use state feedback and do not condition on the predicted actions. We discuss the important architecture decisions in Appendix D.
@@ 3.1 Goal Conditioning
For some tasks such as tracking, a near-future goal can be appended to the input to condition the resulting policy. The conditioning is formalized by defining a set of goal states $\mathcal{G}\subseteq\mathcal{S}$. We then learn a goal-conditioned policy $\pi_{\theta}\bigl(a_{t}\,\big|\;s_{t-1},\dots,s_{t-h^{s}},g\bigr)$ where $g\subseteq\mathcal{G}$. In practice, we define $g$ as a sequence of future states to be reached, i.e., $g=\{s_{i}\}_{i=t}^{t+h^{g}}$, where the goal states are typically generated from a high-level planner that plans instrumental goal states as done in Gupta et al. [11]. In our experiments, we simulate goal states from the demonstration dataset.
@@ 3.2 Action Sampling from a $k$-GMM
QFAT samples the next action from the conditional distribution, $a_{t}\sim\pi_{\theta}(\cdot|s_{t-1},\dots,s_{t-h^{s}},g)$, which is a GMM with $k$ components. In practice, the model can be trained to output a short sequence of actions that are then executed in an open-loop fashion. This is done to ensure action consistency and avoid over-fitting to idle actions [7]. For simplicity, we describe only the next-action generation setting.
@@ 3.3 Stabilizing Output Trajectories
When controlling real-world robotics systems, it is important to generate action sequences without jitter, as high frequency inputs may damage actuators and excite undesirable structural modes [2]. Thanks to the richness of the GMM representation of the policy, we can directly design stabilization patches.
@@
A common heuristic to mitigate noisy samples is to adjust the sampling temperature, which in our GMM parametrization is equivalent to down-scaling the variance in each Gaussian component [29, 16, 50, 51]. Concretely, one replaces each variance $\boldsymbol{\sigma}_{i}^{2}$ by $\alpha\,\boldsymbol{\sigma}_{i}^{2}$, for some small $\alpha\in(0,1)$—sometimes by several orders of magnitude ($\alpha\ll 1$).
This compresses each component, forcing sampled actions to concentrate more tightly around each component’s mean. However, merely shrinking the variances does not address the large-scale spread when mixture components are far apart. The inter-component variance remains, and noisy samples may still arise from small but non-negligible transitions among disparate components.
This phenomenon is exacerbated when the choice of the number of components $k$ is overspecified (see Figure 2 and further discussion in Appendix A). Furthermore, it has been shown by Carreira-Perpinan [5] that GMMs in dimension $d\geq 2$ may have a mode that does not directly align with the mean of any component. In such a situation, the down-scaling technique is likely to result in an insidious loss of the extra mode (see Figure 2).
Mode-tracking via the Mean-Shift Algorithm We explored a principled way to reduce the variance of sampling from multimodal distributions by explicitly identifying the major modes and then sampling from those modes with an appropriate probability while controlling variance. An effective way to find the modes in a GMM is via the mean-shift algorithm [5], which is an efficient fixed-point iteration algorithm (see Appendix B for further details). We propose a multinomial distribution over the modes [^1] The number of modes $J$ is hard to control theoretically [5]. $\{m_{1},...m_{J}\}$ that respects the shape of the density:
@@
$\text{where }w_{j}$ $=p_{\text{GMM}}\bigl(m_{j}\bigr)\cdot\bigl|-H_{j}\bigr|^{-\frac{1}{2}}$ (4)
$\text{and }H_{j}$ $=\nabla^{2}\log p_{\text{GMM}}(m_{j})\,.$ (5)
@@
Figure 2: The figures demonstrate the potential detrimental effects of down-scaling the component variances in a learned Gaussian Mixture Model (GMM). The left figure illustrates how down-scaling variance introduces irreducible variance proportional to the distance between component means, when the number of components ($k$) is misspecified (e.g., approximating a unimodal distribution with two components). The two figures on the right show how reducing variance can lead to a loss of multimodality, specifically causing the central mode to disappear in a 2D two-mixture Gaussian that presents three modes.
@@ 4 Experiments
We evaluate the performance of Q-FAT through extensive experiments across 9 different tasks using five representative simulated robotics environments: PushT [7], Kitchen [11], UR3 BlockPush, BlockPush [10] and Multimodal Ant [26]. These environments cover the most common control signals in practice, namely position and velocity of robot’s end effector and robot joint angles. A detailed description of each environment is provided in Appendix C. We further ran preliminary experiments on an autonomous driving dataset (nuScene [4]) and we add the experimental results in Appendix E. Our experiments are designed to address the following key research questions:
@@
  5. 5.
    How does the inference speed of Q-FAT compare to VQ-BeT?
We find that Q-FAT achieves state-of-the-art performance across the majority of conditional and unconditional tasks, achieving state-of-the-art balance between the quality and diversity of the generated actions. We report the results of Q-FAT using the variance reduced sampling (Section 3.2) with a factor of $10^{-6}$ and we discuss the implications of this choice in Section 4.3. Further training details are provided in Appendix D. A summary of our results is presented in Table 1.